Should print UTKAST on the first two pages, light. Should print UTKAST on bottom of the page after the second page.

\usepackage[light,firsttwo,bottomafter]{draftcopy}

Test7: This is a test page 1

This is a testpage 1 Test7: This is a test page 2 This is a testpage 2 Test7: This is a test page 3 This is a testpage 3 Test7: This is a test page 4 This is a testpage 4